OpenStudy (anonymous):

Find the value of r so that the slope between the points is 4. (6, -2) and (r, -6) a. -5 b. 5 c. -42 d. 4

OpenStudy (anonymous):

Please help!!! I am doing some extra credit homework. I really need to get this correct!!! Thank you.

jigglypuff314 (jigglypuff314):

Hello @CMFink00 and Welcome to OpenStudy! :) \[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]when given points (x1, y1) and (x2, y2) so you can get \[\frac{ -6 + (-2) }{ r - 6 } = 4\]do you know how to solve for r from there? :)

OpenStudy (anonymous):

Um... You, add the top and and you get (-8) then you put -8 as R??? I am confused sorry! I am not good at this...

jigglypuff314 (jigglypuff314):

that's ok ^_^ the -8 is correct but we're not finished there -8 --- = 4 r multiply r on both sides to get r out of the denominator so -8 = 4r then divide both sides by 4 to get r alone

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

-8/4 = -2? and 4/4 =0 so, -2=r?

jigglypuff314 (jigglypuff314):

oh wait! sorry! I forgot that it was a (r - 6) in the denominator >,< so -8 ------- = 4 multiply both sides by (r - 6) (r - 6) and get -8 = 4(r - 6) distribute -8 = 4r - 24 sorry about that, now can you solve for r from there?

OpenStudy (anonymous):

divide -8, 4r, and 24 by 4 right? so -8/4= -2 4r/4=r 24/4=6 right?

OpenStudy (anonymous):

-2=r-6?

jigglypuff314 (jigglypuff314):

yes :) then add 6 to both sides ^_^

OpenStudy (anonymous):

okay so -2+6=4 and -6+6=0 so r=4? Am I correct?

jigglypuff314 (jigglypuff314):

yes! that's what I got :)

OpenStudy (anonymous):

Okay thank you so much!!!

jigglypuff314 (jigglypuff314):

you're welcome :)